# Emergent morphogenesis via planar fabrication enabled by a reduced model of composites

### v1

@@ II.4 Structural similarity index measure
To compare the scanned experimental shapes with numerically simulated ones quantitatively, we use the structural similarity index measure (SSIM), defined at each voxel in a 3D domain. For completeness, we briefly describe our application. We discretize the 3D domain into an $N_{s}\times N_{s}\times N_{s}$ grid and calculate SSIM at each voxel as nilsson2020
@@
where $N_{s}=10$. The voxel intensities are normalized to $0,L_{s}]$, and we use the default parameter values $k_{1}=0.01$, $k_{2}=0.03$, $L_{s}=1$, and $c_{i}=(k_{i}*L_{s})^{2}$ for $i=1,2$ nilsson2020 . Here, $\mu_{1}$, $\mu_{2}$ are local means, $\sigma_{1}^{2}$, $\sigma_{2}^{2}$ the variances, and $\sigma_{12}$ the covariance. To make the scanned and simulated shapes comparable, we uniformly scale, translate, and rotate the simulated shapes to align with the scanned experimental data. For our results, the SSIM ranges from 0 to 1, with 0 indicating no structural correlation and 1 indicating perfect structural similarity.
@@
In Figure 8D, solid markers with error bars indicating standard deviations show the evolution of the aspect ratio $h/d$ over time during thermal actuation. The entire morphing process completes in approximately 15 seconds, dramatically faster than conventional 3D-shape fabrication methods, which often require minutes to hours. Each pattern exhibits a slightly different morphing timeline. The six-arm star on the circular sheet (pattern A) begins deforming earliest at around 5s, and its $h/d$ plateaus after 12s. The two-star pattern on the rectangular sheet (pattern B) begins to morph latest at around 12s. The cross pattern on the square sheet (pattern C) starts at 7s, and ultimately reaches an aspect ratio of $h/d>1.2$. Hollow markers denote the numerical simulations obtained by prescribing thermal strains as described in Sections II.2.2 and II.3. Figure 8E plots the mean SSIM values for the nine SSIM contour plots shown in panels (ii)-(iv). The mean SSIM is simply the arithmetic mean $\overline{\mathrm{SSIM}}=\frac{1}{N_{s}^{3}}\sum_{i=1}^{N_{s}}\sum_{j=1}^{N_{s}}\sum_{k=1}^{N_{s}}\mathrm{SSIM}(i,j,k)$. All mean SSIM values are around 0.8, indicating close agreement between the numerical simulations and the scanned experimental shapes.
@@
III Conclusion
@@
We also developed a robust fabrication protocol to transform planar composites into 3D structures via thermal actuation, achieving a repeatable success rate exceeding 90%. Three representative geometries – a six-arm star on a circle, a two-star on a rectangle, and a cross on a square – showed excellent agreement between numerical simulations and experiments across various morphing stages, as evidenced by both qualitative comparisons in Figures 1, 5-7, and quantitative SSIM values in Figure 8.
@@
IV Experimental Section
Materials and Methods: We used commercially available Shrinky Dink polystyrene sheets (Cridoz Shrinky Art Paper Sheets) and 3D-printed PLA kirigami inspired patterns. The kirigami components were designed using Fusion 360 and SolidWorks CAD software solidworks , utilizing Python scripts to rapidly assign and vary geometry. The designs were fabricated with a fused deposition modelling 3D printer (Bambu Lab 1PS) using PLA filament chosen for its thermal resistance during oven experimentation and its mechanical properties. Adhesion between layers was achieved using a two-part Gorilla epoxy adhesive, selected for its bond strength, thermal stability, cost, and ease of use.
Shrinky Dink Properties: Shrinky Dink is a thermoplastic polystyrene polymer that exhibits isotropic in-plane shrinkage when heated. We define the shrink ratio as $L/L_{0}$, where $L$ is the instantaneous length and $L_{0}$ is the original length. Figure 3 shows the variation of normalized length $\bigl{(}L/L_{0}\bigr{)}$ with normalized temperature $T/T_{g}$. The shrink ratio decreases with increasing temperature and plateaus once $T/T_{g}$ exceeds 1.15, indicating isotropic behavior as the shrink ratios along orthogonal in-plane directions converge. Above $300\,^{\circ}\mathrm{F}$, the $x$- and $y$-axis gauge lengths coincide and the standard deviation bars contract, reflecting highly reproducible shrinkage. To ensure uniform and predictable actuation, all subsequent experiments were conducted within this temperature regime.
@@ IV Experimental Section
    Cure under distributed weight $(6.8\times 10^{2}\ \mathrm{N/m^{2}})$ for 24 h at room temperature to ensure full adhesion.
@@
Pattern Design and Results: Morphing samples were generated via a Python script that produced random polygonal kirigami patterns on a 2D domain. Visual inspection of the resulting 3D shapes revealed that local curvature propagated along the printed line directions. In a parametric study, we varied three parameters: line width, PLA layer thickness, and line count per polygon. Empirically, increasing line count yielded smoother and more continuous circumferential curvature, whereas fewer lines produced segmented outlines. Thicker and wider PLA lines resisted actuation, requiring longer oven exposure due to higher bending and stretching stiffness. Consequently, designs with more than four lines were adopted for rapid, reliable thermal response and consistent circumferential curvature across samples. Three planar patterns were therefore designed to benchmark the aforementioned observations and to investigate the numerical accuracy of simulations in comparison to the experimentally fabricated patterns. They are: A. Single-star: Fabricated by intersecting three coplanar “arms”, each $90\,\mathrm{mm}$ long, at $60^{\circ}$ increments through the centroid, generating six $45\,\mathrm{mm}$ radial arms. B. Double-star: Superimposed two such stars aligned coaxially, separated by a $56.73\,\mathrm{mm}$ spine, giving an overall tip-to-tip length of $146.7\,\mathrm{mm}$. C. Orthogonal X-cross: Incorporating four arms spaced at $90^{\circ}$ intervals, each extending $56.7\,\mathrm{mm}$ from the center of the shape and spanning $99.99\,\mathrm{mm}$ along the principal diagonal, with each arm chamfered at its end via the addition of a right-angled triangle.
@@
Computational Reconstruction: The computational reconstruction utilized Polycam photogrammetry software alongside a high-resolution camera and a custom-printed stand to capture 80 images around each sample from multiple angles. These images were processed in Polycam to generate and edit 3D models in STL and OBJ formats. To enhance reconstruction consistency and spatial accuracy, we applied fiduciary markers to the samples prior to imaging, facilitating reliable feature recognition and reducing computation time surmen2023 . The resulting meshes were scaled to real-world dimensions based on known physical measurements and subsequently 3D-printed to validate scale fidelity against experimental results.
@@
Aspect Ratio: To extract the mean aspect ratio of each experimental sample, we conducted a frame-by-frame analysis at 1 s intervals using a high resolution video of the thermal actuation procedure. Frames were imported into ImageJ imagej2025 , where a threshold mask isolated the sample from the background for accurate perimeter and geometry detection. The straight-line tool was used to manually define key dimensions, whose pixel lengths were recorded. Conversion from pixels to physical units was performed using a known reference length inside the oven. From these measurements, in-plane shrinkage and out-of-plane height expansion were computed and expressed as the numerical aspect ratio $h/d$, as demonstrated in Figure 8.
@@
V Appendix A. Gradient and Hessian of coupled Energy
We consider axial strain caused by temperature change or prestretch. The thermal strain is given by
$\varepsilon^{\mathrm{th}}=\alpha\,(T-T_{0})$
The gradient and Hessian of the new energy formulation, Equation 6, are given by
$\nabla_{\bm{X}}E(\bm{X})$ $=\sum_{i=0}^{N_{e}-1}k_{s}\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bigl{[}\varepsilon_{i}(\bm{X})-\varepsilon_{i}^{\mathrm{th}}\bigr{]}}\,\nabla_{\bm{X}}\varepsilon_{i}(\bm{X})$ (12)
$\quad+\sum_{j=0}^{N_{h}-1}k_{b}\,\bigl{[}\theta_{j}(\bm{X})-\beta l_{0}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bigl{(}\Delta\varepsilon_{j}(\bm{X})-\Delta\varepsilon_{j}^{\mathrm{th}}\bigr{)}}\bigr{]}\,\Bigl{[}\nabla_{\bm{X}}\theta_{j}(\bm{X})-\beta l_{0}\sum_{p\in\partial j}s_{p}\,\nabla_{\bm{X}}\varepsilon_{jp}(\bm{X})\Bigr{]},$
and
$\nabla_{\bm{X}}^{2}E(\bm{X})$ $=\sum_{i=0}^{N_{e}-1}k_{s}\Bigl{(}\nabla_{\bm{X}}\varepsilon_{i}\otimes\nabla_{\bm{X}}\varepsilon_{i}+\;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bigl{[}\varepsilon_{i}(\bm{X})-\varepsilon_{i}^{\mathrm{th}}\bigr{]}}\,\nabla_{\bm{X}}^{2}\varepsilon_{i}\Bigr{)}$ (13)
$\quad+\sum_{j=0}^{N_{h}-1}k_{b}\Bigl{(}\bigl{[}\nabla_{\bm{X}}\theta_{j}-\beta l_{0}\sum_{p\in\partial j}s_{p}\,\nabla_{\bm{X}}\varepsilon_{jp}\bigr{]}\otimes\bigl{[}\nabla_{\bm{X}}\theta_{j}-\beta l_{0}\sum_{p\in\partial j}s_{p}\,\nabla_{\bm{X}}\varepsilon_{jp}\bigr{]}$
$\qquad\quad+\,\bigl{[}\theta_{j}(\bm{X})-\beta l_{0}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bigl{(}\Delta\varepsilon_{j}(\bm{X})-\Delta\varepsilon_{j}^{\mathrm{th}}\bigr{)}}\bigr{]}\,\Bigl{(}\nabla_{\bm{X}}^{2}\theta_{j}-\beta l_{0}\sum_{p\in\partial j}s_{p}\,\nabla_{\bm{X}}^{2}\varepsilon_{jp}\Bigr{)}\Bigr{)}\!.$
Finite-difference approximations have been used to verify the formulas.
VI Appendix B. Dimensions of bilayer composites

### v2

@@ 2.4 Structural Similarity Index Measure
To compare the scanned experimental shapes with numerically simulated ones quantitatively, we use the structural similarity index measure (SSIM), defined at each voxel in a 3D domain. For completeness, we briefly describe our application. We discretize the 3D domain into an $N_{s}\times N_{s}\times N_{s}$ grid and calculate SSIM at each voxel as [36]
@@
where $N_{s}=10$. The voxel intensities are normalized to $0,L_{s}]$, and we use the default parameter values $k_{1}=0.01$, $k_{2}=0.03$, $L_{s}=1$, and $c_{i}=(k_{i}*L_{s})^{2}$ for $i=1,2$ [36]. Here, $\mu_{1}$, $\mu_{2}$ are local means, $\sigma_{1}^{2}$, $\sigma_{2}^{2}$ the variances, and $\sigma_{12}$ the covariance. To make the scanned and simulated shapes comparable, we uniformly scale, translate, and rotate the simulated shapes to align with the scanned experimental data. For our results, the SSIM ranges from 0 to 1, with 0 indicating no structural correlation and 1 indicating perfect structural similarity.
@@
In Figure 8D, solid markers with error bars indicating standard deviations show the evolution of the aspect ratio $h/d$ over time during thermal actuation. The entire morphing process completes in approximately 15 seconds, dramatically faster than conventional 3D-shape fabrication methods, which often require minutes to hours. Each pattern exhibits a slightly different morphing timeline. The six-arm star on the circular sheet (pattern A) begins deforming earliest at around 5s, and its $h/d$ plateaus after 12s. The two-star pattern on the rectangular sheet (pattern B) begins to morph latest at around 12s. The cross pattern on the square sheet (pattern C) starts at 7s, and ultimately reaches an aspect ratio of $h/d>1.2$. Hollow markers denote the numerical simulations obtained by prescribing thermal strains as described in Sections 2.2.2 and 2.3. Figure 8E plots the mean SSIM values for the nine SSIM contour plots shown in panels (ii)-(iv). The mean SSIM is given by $\overline{\mathrm{SSIM}}=\tfrac{1}{N_{s}^{3}}\sum_{i=1}^{N_{s}}\sum_{j=1}^{N_{s}}\sum_{k=1}^{N_{s}}\mathrm{SSIM}(i,j,k)$. All mean SSIM values are around 0.8, indicating close agreement between the numerical simulations and the scanned experimental shapes.
The restriction on the choice of actuation duration (16 s, 15 s, and 12 s for patterns A to C, respectively) is that it must fall within an appropriate range. In experiments, significantly longer heating times would lead to excessive material softening or partial melting of the composite, resulting in uncontrolled deformation. For the materials used in this study, thermal actuation durations within approximately 20 s remain mechanically stable. Accordingly, we select actuation durations that yield visually representative morphing configurations and clearly demonstrate the design capability, while avoiding excessive softening. With longer heating times, both the Shrinky Dink substrate and the PLA components soften significantly. Due to gravity and the reduced stretching and bending stiffnesses, the structures may fold inward or collapse, transitioning beyond the elastic deformation regime considered in our model. Since the present model focuses on elastic thermo-mechanical coupling, samples are removed from the oven before significant softening or melting occurs.
Importantly, the objective of the thermal actuation is not to reach a steady state. The horizontal axis in Figure 8D represents the prescribed actuation duration rather than physical time in a dynamic sense. Thermal actuation can be readily controlled by removing the sample from the oven and allowing it to cool to room temperature, at which point deformation effectively ceases. As a result, reaching a steady state is not required for either the experiments or the simulations.
The different evolution rates of the aspect ratio $h/d$ among patterns arise partially from geometric effects. Composites with patterns B and C contain sharper corners than pattern A, which contribute more noticeably to out-of-plane deformation during actuation. Consequently, the aspect ratio in patterns B and C are more sensitive to out-of-plane deformation than in pattern A under the same thermal actuation process.
@@
2.5 Numerical Results Verified with Finite Element Simulations
We further verify the results from discrete elastic shell (DES) model by comparison with finite elements (FE) simulations performed using Abaqus [16]. The bilayer composite is modeled as two shell layers, each discretized with S3R shell elements. Each layer contains $N$ nodes, resulting in a total of $2N$ nodes. The two layers are coupled through a tie constraint between the contacting surfaces. Both layers are modeled as isotropic Mooney-Rivlin hyperelastic materials with strain energy density [16]
$W(\mathbf{F})=C_{1}(\bar{I}_{1}-3)+C_{2}(\bar{I}_{2}-3)+\frac{1}{D_{1}}(J-1)^{2},$ (26)
where $\bar{I}_{1}$ and $\bar{I}_{2}$ are the first and second invariants of the isochoric right Cauchy-Green tensor, respectively, $J$ is the determinant of the deformation gradient, and $C_{1}$, $C_{2}$ and $D_{1}$ are fitting parameters for materials.
In the infinitesimal strain limit, the Mooney-Rivlin model can be linearized to an equivalent isotropic elastic solid characterized by shear modulus $G$ and bulk modulus $K$ with
$C_{1}=C_{2}=\frac{G}{4},\qquad D_{1}=\frac{2}{K},$ (27)
where the derivation is provided in Appendix 6 under the assumption $C_{1}=C_{2}$.
Recall that during thermal actuation, the Young’s moduli are taken as $Y_{0,1}=0.8911$ MPa for the Shrinky Dink substrate layer and $Y_{0,2}=2.6733$ MPa for the PLA kirigami layer, reduced from their room temperature values as discussed in Section 2.2. The Poisson’s ratio is $0.33$ for both materials. This results in $C1=C2=8.38\times 10^{4}$Pa and $D_{1}=2.29\times 10^{-6}\rm{Pa}^{-1}$ for Shrinky Dink polystyrene, and $C1=C2=2.51\times 10^{5}$Pa and $D_{1}=7.63\times 10^{-7}\rm{Pa}^{-1}$ for the PLA layer.
We modeled the deformation using an explicit time integration in Abaqus [16]. The mass density is set to $1.05\rm{g/cm}^{3}$ for the Shrinky Dink Polystyrene substrate layer [5], and $1.25\rm{g/cm}^{3}$ for the kirigami PLA layer [28]. The thickness of the Shrinky Dink layer is 0.3mm. The thickness of PLA layer is 0.7mm for pattern A, 0.6 mm for pattern B, and 1.0mm for pattern C. The kirigami layer has a zero thermal expansion coefficient, whereas the substrate layer has a nonzero thermal expansion coefficient. Each node in the FE model has 6 DOFs.
Figure 9 shows the verification of the DES model against FE simulations for two representative cases. The first row (A) corresponds to pattern A in Figure 4 with thermal strain parameter $\varepsilon^{\rm th}_{\rm pre}=-0.77$, and the second row (B) corresponds to pattern C with $\varepsilon^{\rm th}_{\rm pre}=-0.50$. The first column shows the deformed configuration, and the second column shows the contours of the minimum principal logarithmic strain $\varepsilon^{\rm LE}_{\min}$, the FE simulations closely reproduce the experimentally scanned configurations and show good agreement with DES results in Figures 5 and 7. The last column shows the top-view SSIM contours computed between the FE and DES results. The mean SSIM values are 0.83 and 0.81 for cases A(iii) and B(iii), respectively, demonstrating quantitative agreement between the DES and FE simulations.
Figure 9: Verification of discrete elastic shell (DES) simulations against finite element (FE) results using Abaqus. The first row (A) shows results for pattern A with thermal strain parameter $\varepsilon^{\rm th}_{\rm pre}=-0.77$, and the second row (B) shows the results for pattern C with $\varepsilon^{\rm th}_{\rm pre}=-0.50$. Column (i) shows the deformed configurations, column (ii) shows contours of the minimum principal logarithmic strain $\varepsilon^{\rm LE}_{\min}$, and column (iii) shows the top-view SSIM contours computed between the FE and DES results. The mean SSIM values are 0.83 and 0.81 for A(iii) and B(iii), respectively.
For the bilayer composite considered here, the FE model represents the structure as two shell layers coupled by a tie constraint. Each layer contains N nodes connected by S3R shell elements, resulting in a total of 2N nodes with six degrees of freedom per node. Both layers are modeled as Mooney-Rivlin hyperelastic materials. In contrast, the DES method models the bilayer composite as a reduced single layer shell with N nodes and 3 DOFs per node, using the elastic energy in Equation 6. Using the same numerical mesh and temperature change (i.e., thermal strain) distribution, the CPU time required to complete the simulation is 738 s for the FE model and 420 s for the DES model for pattern A, and 218 s for the FE model and 183 s for the DES model for pattern C, on a personal computer [^1] A Dell Inspiron 3891 with an Intel Core i7-10700 CPU (8 cores) at 2.90 GHz processor, 12 GB of memory, a 512 GB SSD, and a 2 TB HDD (7200 RPM).. The improved computational efficiency of the DES approach arises from the reduced number of DOFs and numerical acceleration techniques, including sparse matrix assembly and sparse LU factorization (via the splu function in SciPy) at each Newton iteration.
3 Conclusion
@@
We also developed a robust fabrication protocol to transform planar composites into 3D structures via thermal actuation, achieving a repeatable success rate exceeding 90%. Three representative geometries – a six-arm star on a circle, a two-star on a rectangle, and a cross on a square – showed excellent agreement between numerical simulations and experiments across various morphing stages, as evidenced by both qualitative comparisons in Figures 5-7, and quantitative SSIM values in Figure 8.
@@
4 Experimental Section
Materials and Methods: We used commercially available Shrinky Dink polystyrene sheets (Cridoz Shrinky Art Paper Sheets) and 3D-printed PLA kirigami inspired patterns. The kirigami components were designed using Fusion 360 and SolidWorks CAD software [17], utilizing Python scripts to rapidly assign and vary geometry. The designs were fabricated with a fused deposition modelling 3D printer (Bambu Lab 1PS) using PLA filament chosen for its thermal resistance during oven experimentation and its mechanical properties. Adhesion between layers was achieved using a two-part Gorilla epoxy adhesive, selected for its bond strength, thermal stability, cost, and ease of use.
Shrinky Dink Properties: Shrinky Dink is a thermoplastic polystyrene polymer that exhibits isotropic in-plane shrinkage when heated. We define the shrink ratio as $L/L_{0}$, where $L$ is the instantaneous length and $L_{0}$ is the original length. Figure 3E shows the variation of normalized length $\bigl(L/L_{0}\bigr)$ with normalized temperature $T/T_{g}$. The shrink ratio decreases with increasing temperature and plateaus once $T/T_{g}$ exceeds 1.15, indicating isotropic behavior as the shrink ratios along orthogonal in-plane directions are similar. Above $300\,^{\circ}\mathrm{F}$, the $x$- and $y$-axis gauge lengths coincide and the standard deviation bars contract, reflecting highly reproducible shrinkage. To ensure uniform and predictable actuation, all subsequent experiments were conducted within this temperature regime.
@@ 4 Experimental Section
    Cure under distributed weight $(6.8\times 10^{2}\ \mathrm{N/m^{2}})$ over 10 h at room temperature to ensure full adhesion.
@@
Pattern Design and Results: Morphing samples were generated via a Python script that produced random polygonal kirigami patterns on a 2D domain. Visual inspection of the resulting 3D shapes revealed that local curvature propagated along the printed line directions. In a parametric study, we varied three parameters: line width, PLA layer thickness, and line count per polygon. Empirically, increasing line count yielded smoother and more continuous circumferential curvature, whereas fewer lines produced segmented outlines. Thicker and wider PLA lines resisted actuation, requiring longer oven exposure due to higher bending and stretching stiffness. Consequently, designs with more than four lines were adopted for rapid, reliable thermal response and consistent circumferential curvature across samples. Three planar patterns were therefore designed to benchmark the aforementioned observations and to investigate the numerical accuracy of simulations in comparison to the experimentally fabricated patterns. They are: A. Single-star: Fabricated by intersecting three coplanar “arms”, each $90\,\mathrm{mm}$ long, at $60^{\circ}$ increments through the centroid, generating six $45\,\mathrm{mm}$ radial arms. B. Double-star: Superimposed two such stars aligned coaxially, separated by a $56.73\,\mathrm{mm}$ spine, giving an overall tip-to-tip length of $146.7\,\mathrm{mm}$. C. Orthogonal X-cross: Incorporating four arms spaced at $90^{\circ}$ intervals, each extending $56.7\,\mathrm{mm}$ from the center of the shape and spanning $99.99\,\mathrm{mm}$ along the principal diagonal, with each arm chamfered at its end via the addition of a right-angled triangle, Figure 9.
@@
Computational Reconstruction: The computational reconstruction utilized Polycam photogrammetry software alongside a high-resolution camera and a custom-printed stand to capture 80 images around each sample from multiple angles. These images were processed in Polycam to generate and edit 3D models in STL and OBJ formats. To enhance reconstruction consistency and spatial accuracy, we applied fiduciary markers to the samples prior to imaging, facilitating reliable feature recognition and reducing computation time [45]. The resulting meshes were scaled to real-world dimensions based on known physical measurements and subsequently 3D-printed to validate scale fidelity against experimental results.
@@
Aspect Ratio: To extract the mean aspect ratio of each experimental sample, we conducted a frame-by-frame analysis at 1 s intervals using a high resolution video of the thermal actuation procedure. Frames were imported into ImageJ [37], where a threshold mask isolated the sample from the background for accurate perimeter and geometry detection. The straight-line tool was used to manually define key dimensions, whose pixel lengths were recorded. Conversion from pixels to physical units was performed using a known reference length inside the oven. From these measurements, in-plane shrinkage and out-of-plane height expansion were computed and expressed as the numerical aspect ratio $h/d$, as demonstrated in Figure 8.
@@
5 Appendix A. Asymptotic Analysis of Thermal Strain Distribution
We assume an exponentially varying thermal strain field
$\varepsilon^{\mathrm{th}}(d)=\varepsilon_{\mathrm{pre}}^{\mathrm{th}}\,\frac{1-\exp\!\left(-{d}/{L_{c}}\right)}{1-\exp\!\left(-{d_{\max}}/{L_{c}}\right)},$ (28)
where $d\in[0,d_{\max}]$ denotes the distance from an edge midpoint to the closet kirigami pattern boundary, $L_{c}>0$ is a characteristic decay length controlling the spatial localization, and $\varepsilon_{\mathrm{pre}}^{\mathrm{th}}$ is the prescribed thermal strain parameter. This ensures $\varepsilon^{\mathrm{th}}(0)=0$ and $\varepsilon^{\mathrm{th}}(d_{\max})=\varepsilon_{\mathrm{pre}}^{\mathrm{th}}$.
We define a dimensionless localization parameter $\gamma\in(0,\infty)$ to relate $L_{c}$ with $d_{\max}$
$L_{c}={d_{\max}}/{\gamma},$ (29)
Equation (28) can be expressed in nondimensional form as
$\varepsilon^{\mathrm{th}}(d/d_{\max})=\varepsilon_{\mathrm{pre}}^{\mathrm{th}}\,\frac{1-\exp(-\gamma d/d_{\max})}{1-\exp(-\gamma)}.$ (30)
In the limit $\gamma\to 0$, a Taylor expansion of the exponential term is
$\exp(-\gamma d/d_{\max})=1-\gamma d/d_{\max}+O(\gamma^{2}).$ (31)
Then Equation (30) reduces to
$\lim_{\gamma\to 0}\varepsilon^{\mathrm{th}}(\frac{d}{d_{\max}})=\varepsilon_{\mathrm{pre}}^{\mathrm{th}}\frac{d}{d_{\max}},$ (32)
which corresponds to a linear thermal strain distribution.
On the other hand, in the limit $\gamma\to\infty$, Equation (30) reduces to
$\lim_{\gamma\to\infty}\varepsilon^{\mathrm{th}}(\frac{d}{d_{\max}})=\varepsilon_{\mathrm{pre}}^{\mathrm{th}},\quad d>0,$ (33)
while $\varepsilon^{\mathrm{th}}(0)=0$ for all $\gamma$. This limit corresponds to a strongly localized thermal strain near the kirigami pattern boundary.
6 Appendix B. Equivalent Mooney-Rivlin Parameters from Infinitesimal Strain Linearization
For linear elasticity, we have $\sigma_{ij}=C_{ijkl}\,\varepsilon_{kl}$. For an isotropic material, this relation can be written in Lamé form as
$\boldsymbol{\sigma}=2\mu\,\boldsymbol{\varepsilon}+\lambda(\operatorname{tr}\boldsymbol{\varepsilon})\mathbf{I},$
where $\mu$ and $\lambda$ are the Lamé parameters. Equivalently, the Cauchy stress can be expressed as
$\boldsymbol{\sigma}=2\mu\,\boldsymbol{\varepsilon}^{\prime}+\left(\lambda+\frac{2\mu}{3}\right)(\operatorname{tr}\boldsymbol{\varepsilon})\mathbf{I}=2G\,\boldsymbol{\varepsilon}^{\prime}+K(\operatorname{tr}\boldsymbol{\varepsilon})\mathbf{I},$ (34)
where, the deviatoric strain $\boldsymbol{\varepsilon}^{\prime}=\boldsymbol{\varepsilon}-\frac{1}{3}(\operatorname{tr}\boldsymbol{\varepsilon})\mathbf{I}$, shear modulus $G=\mu$, and bulk modulus $K=\lambda+\frac{2}{3}\mu$.
With Equation 34, the linear elastic energy density is given by
$W_{\mathrm{lin}}=\frac{1}{2}\,\boldsymbol{\sigma}:\boldsymbol{\varepsilon}=\frac{1}{2}\Big(2G\,\boldsymbol{\varepsilon}^{\prime}+K(\operatorname{tr}\boldsymbol{\varepsilon})\mathbf{I}\Big):\boldsymbol{\varepsilon}=G\,\boldsymbol{\varepsilon}^{\prime}:\boldsymbol{\varepsilon}^{\prime}+\frac{K}{2}(\operatorname{tr}\boldsymbol{\varepsilon})^{2},$ (35)
where Equation 34 was substituted, and $\boldsymbol{\varepsilon}^{\prime}:\mathbf{I}=0$ and $\mathbf{I}:\boldsymbol{\varepsilon}=\operatorname{tr}\boldsymbol{\varepsilon}$ were used.
For an isotropic, compressible hyperelastic material, the Mooney-Rivlin strain energy density is given by [16]
$W(\mathbf{F})=C_{1}(\bar{I}_{1}-3)+C_{2}(\bar{I}_{2}-3)+\frac{1}{D_{1}}(J-1)^{2},$ (36)
where
$\bar{I}_{1}=J^{-2/3}I_{1},\quad\bar{I}_{2}=J^{-4/3}I_{2},\quad I_{1}=\mathrm{tr}\,\mathbf{C},\quad I_{2}=\tfrac{1}{2}[(\mathrm{tr}\,\mathbf{C})^{2}-\mathrm{tr}(\mathbf{C}^{2})],\quad J=\det\mathbf{F},\quad\mathbf{C}=\mathbf{F}^{T}\mathbf{F}.$
The infinitesimal strain limit of Equation 36 gives [1, 26]
$W=2(C_{1}+C_{2})\,\varepsilon^{\prime}:\varepsilon^{\prime}+\frac{1}{D_{1}}\left(\mathrm{tr}\,\varepsilon\right)^{2}.$ (37)
Comparing Equations 35 and 37, the equivalent shear and bulk moduli are given by
$G=2(C_{1}+C_{2}),\qquad K=2/D_{1}.$ (38)
Assuming $C_{1}=C_{2}$, we obtain $C_{1}=C_{2}=G/4$.
7 Appendix C. Dimensions of Bilayer Composites